OpenStudy (anonymous):

reparametarize the curve "alpha" in alpha(t)=(e^t,e^-t,2t) using h(s)=logs on j:s>o

OpenStudy (zehanz):

Do you mean t=logs?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

yes

OpenStudy (zehanz):

OK, so \(\alpha(s)=(e^{\log s}, e^{-\log s}, 2\log s) \). I think that you also mean that \(\log s= \log_{e}s=\ln s\)...if so, you can simplify easily.

OpenStudy (anonymous):

what is the answer?